Use the Cauchy-Schwarz inequality to solve x³-25x²-4x+100=0, if we know
that all its roots are rational.
1
Expert's answer
2020-02-25T08:30:41-0500

As per the given question,

"x^3-25x^2-4x+100=0,"

Now,

"x^3-4x-25x^2+100=0,"

"x(x^2-4)-25(x^2-4)=0,"

"(x^2-4)(x-25)=0,"

so,

"x^2=4" or "x=25",

"x=\pm2" or "x=25".